How do you solve this?
Find the equation of the line passing through (3,-1) and perpendicular to the line 3𝑥 +9 𝑦 = −2.

Answer by Alan3354(69443)

How do you solve this?
Find the equation of the line passing through (3,-1) and perpendicular to the line 3𝑥 +9 𝑦 = −2.
---------------
Step 1, find the slope of the given line.
Step 2, the slope of lines perpendicular to it is the negative inverse. Call it m.
Step 3, use y-y1 = m*(x-x1) where (x1,y1) is the given point.